OpenStudy (anonymous):

Simplify the expression. A. 0 B. 0.5 C. 1 D. 50

OpenStudy (anonymous):



OpenStudy (charlotte123):

so its just 50? :P then it would be 50 :D

OpenStudy (anonymous):

any number raised to the 0 power = 1. so 5^0 = 1; and (345)^0 = 1, and (-2x3)^0 = 1

OpenStudy (arbershabani97):

it is for example